Show that 2z=(ax+y)^2+b,where a,b are arbitrary constants is a complete
integral of px+qy-q^2=0
1
Expert's answer
2020-03-29T09:48:05-0400
"2z=(ax+y)^2+b\rightarrow\boxed{z(x,y)=\frac{1}{2}(ax+y)^2+\frac{b}{2}}"



Then,



"p=\frac{\partial z}{\partial x}=\frac{1}{2}\cdot 2\cdot(ax+y)\cdot a=a^2x+ay\\[0.3cm]
q=\frac{\partial z}{\partial y}=\frac{1}{2}\cdot 2\cdot(ax+y)=ax+y\\[0.3cm]
px+qy-q^2=(a^2x+ay)x+(ax+y)y-(ax+y)^2=\\[0.3cm]
=a^2x^2+axy+axy+y^2-(a^2x^2+2axy+y^2)=\\[0.3cm]
=a^2x^2+2axy+y^2-a^2x^2-2axy-y^2=0"

Conclusion,



"\boxed{2z=(ax+y)^2+b\rightarrow px+qy-q^2=0}"